A nozzle of inner radius 1.04 mm is connected to a hose of inner radius 9.35 mm. The nozzle shoots out water moving at 25.0 m/s.

What is the mass flow rate?
Answer #1

Mass flow rate = mass of liquid flowing / time

= volume × density / time

= area × length × density / time

= area × speed × density

= × (1.04×10-3)2 × 25 × 1000

= (22/7) × 1.0816 × 25 × 10-3

= 84.98 × 10-3 kg/s

[ Note - Volume flow rate = area × speed

    = 84.98 × 10-6 m3/s ]